OpenStudy (anonymous):

Which functions in the table below give values that could come from exponential functions? Check all that apply. x -3 -2 -1 0 1 2 3 f(x) 0.67 1.11 1.82 3 5.95 9.15 14.45 g(x) 0.125 0.25 0.5 1 2 4 8 h(x) 0.25 0.5 0.75 1 1.25 1.5 1.75 k(x) 64 16 4 1 0.25 0.0625 0.0156

OpenStudy (campbell_st):

g(x) each value is double the previous k(x) each value is 1/4 of the previous...

OpenStudy (anonymous):

so what does that mean?? @campbell_st

OpenStudy (campbell_st):

ok... here is the exponential function for g(x) \[g(x) = 2^x\] so substitute x = -3 you get \[g(-3) = 2^{-3} = 0.125\] for k(x) you get \[k(x) = (\frac{1}{4})^{x}\] so \[k(-3) = (\frac{1}{4})^{-3} = 64\]

OpenStudy (anonymous):

ok i undertand now thank you very much